The area of a rectangle is 32 square inches. The length is 4 inches longer than the width. What is the length? What is the width? What is the perimeter?
Answer by oberobic(2304)

The area of a rectangle = length * width = l * w = 32 sq inches
.
l = w + 4 :: given
.
substituting
(w+4) * w = 32
w^2 + 4w = 32
w^2 + 4w - 32 = 0
.
factoring
(w+8)(w-4) = 0
.
so the solutions are w=-8 and w=4.
of course, a negative length is nonsense, so the only answer appears to be w=4
.
given l = w+4, then l= 4+4 = 8
.
checking the area
4*8 = 32??
yes
.
the perimeter of a rectangle = 2l + 2w = 2(8) + 2(4) = 16+8 = 24
.
Answer:
Length = 8
Width = 4
Perimeter = 24
.
done